OpenStudy (anonymous):

A volleyball team had 6 boys and 5 girls on it. The team needed 2 equipment managers. The coach selected 1 player at random to be an equipment manager and then selected 1 of the remaining players at random to be the other equipment manager. What is the probability that both equipment managers were girls? Express your answer in simplest form.

OpenStudy (anonymous):

5/6 I think

OpenStudy (anonymous):

NO ive had this problem before i just forgot its either 2/11 or 5/11 i just cant figure it out

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

(5/11)*(4/10) = 2/11

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No Problem ;)

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

At first you have 5 girls out of 11 boys/girls on the team. The probability of choosing a girl is 5/11, assuming that we had a success, now we have 4 girls out of 10 boys/girls remaining that can be choosed, now the probability of choosing another girl is 4/10, so... (5/11)*(4/10) = 2/11